OpenStudy (anonymous):

An investor invested a total of $2500 in two mutual funds. One fund earned a profit of 4% and the other earned a 2% profit. If the investor's total profit was $70 how much was invested in to each fund?

OpenStudy (anonymous):

x,y= Investments in each of the firms x+y=2500 (i) also, given 4% of x + 2% of y =70 --> 0.04x + 0.02y=70 (ii) Multiplying (i) by 0.02 0.02x + 0.02y= 50 (iii) Subtracting (iii) from (ii), 0.02x=20 -->x=1000 so y=1500 Answer: $1000 invested in 4% profit fund and $1500 invested in 2% profit fund

OpenStudy (anonymous):

Thank you that makes total sense now that I see it